OpenStudy (anonymous):

solve: 2sin(2x)-2sinx+2√(3) cosx-√(3)=0

OpenStudy (anonymous):

x=\[\pi\]/3

OpenStudy (dumbcow):

x could also be \[4\pi/3 , 5\pi/3\]